### v2

# Predicting O-GlcNAcylation Sites in Mammalian Proteins with Transformers and RNNs Trained with a New Loss Function

Abstract
Glycosylation, a protein modification, has multiple essential functional and structural roles. O-GlcNAcylation, a subtype of glycosylation, has the potential to be an important target for therapeutics, but methods to reliably predict O-GlcNAcylation sites had not been available until 2023; a 2021 review correctly noted that published models were insufficient and failed to generalize. Moreover, many are no longer usable. In 2023, a considerably better recurrent neural network (RNN) model with an F1 score of 36.17% and a Matthews Correlation Coefficient (MCC) of 34.57% on a large dataset was published. This article first sought to improve these metrics using transformer encoders. While transformers displayed high performance on this dataset, their performance was inferior to that of the previously published RNN. We then created a new loss function, which we call the weighted focal differentiable MCC, to improve the performance of classification models. RNN models trained with this new function display superior performance to models trained using the weighted cross-entropy loss; this new function can also be used to fine-tune trained models. A two-cell RNN trained with this loss achieves state-of-the-art performance in O-GlcNAcylation site prediction with an F1 score of 38.88% and an MCC of 38.20% on that same large dataset.
1 Introduction
Glycosylation, a co- and post-translational modification, occurs when glycan(s) are added to proteins. O-linked glycosylation occurs when a glycan is added to an oxygen of an amino acid (usually serine or threonine in mammals). O-GlcNAcylation, a subtype of O-linked glycosylation, occurs when the first glycan added is an N-Acetylglucosamine (abbreviated GlcNAc) (Schjoldager et al., 2020). Unlike other forms of glycosylation, O-GlcNAcylation may be viewed similarly to phosphorylation, for no other glycans are further added to a site after GlcNAc and the process is catalyzed solely by two enzymes, OGT and OGA. O-GlcNAcylation is important functionally and structurally (Schjoldager et al., 2020; Chang et al., 2020); conversely, incorrect O-GlcNAcylation or its improper removal is associated with multiple diseases such as cancers (Shi et al., 2022), infections (Chang et al., 2020), and heart failure (Umapathi et al., 2021). According to recent research, O-GlcNAcylation can be a powerful target for therapeutics (Zhu & Hart, 2021), further emphasizing its relevance.
While O-GlcNAcylation is essential for human health and very important for biotherapeutics, challenges still exist. Because O-GlcNAcylation is complete after the addition of a single glycan, the problem of predicting glycan distributions in a glycosylation site does not exist; thus, the main predictive task is to determine where and when (if ever) an amino acid will be O-GlcNAcylated. As summarized by Seber & Braatz (2024), this task is challenging for “multiple reasons”, which include “a low frequency of events (only about 2% of S/T sites are O-GlcNAcylated) and a lack of a motif to guide predictive efforts”. Moreover, Seber & Braatz (2024) also note that “the effects of neighboring amino acids likely influence whether an S/T is O-GlcNAcylated”, a fact corroborated by their interpretability studies done with their final model, but many previous works use model architectures that cannot directly take into account any type of sequential information. The pioneer classifier machine-learning model for this challenging task is YinOYang (YoY) (Gupta & Brunak, 2002), which also had one of the highest performances until the work of Seber & Braatz (2024), first published in 2023. However, until at least 2021, O-GlcNAcylation site prediction models had insufficient performance to help advance research in the area (Mauri et al., 2021). Specifically, Mauri et al. (2021) found that no published model until then could achieve a precision $\geq$ 9% on a medium-sized independent dataset, suggesting that models to predict the location of O-GlcNAcylation sites failed to generalize successfully in spite of the high metrics achieved with their training data. The models evaluated by Mauri et al. (2021) also have low F1 scores and Matthew Correlation Coefficients (MCCs), another sign that their performance is insufficient. Seber & Braatz (2024) used recurrent neural network (RNN) models (specifically, long short-term memory [LSTM] models) to predict the presence of O-GlcNAc sites from mammalian protein sequence data. That work used two different datasets (from Mauri et al. (2021); Wulff-Fuentes et al. (2021)), the latter of which is among the largest O-GlcNAcylation datasets available. These RNNs obtained vastly superior performance to the previously published models, achieving an F1 score more than 3.5-fold higher and an MCC more than 4.5-fold higher than the previous state of the art. Furthermore, Shapley values were used to interpret the predictions of that RNN model through the sum of simple linear coefficients (Shapley, 1951). These predictions with Shapley values maintained most of the performance of the original RNN models (Supplemental data of Seber & Braatz (2024)). Conversely, two other models that were also released recently, LM-OGlcNAc-Site (Pokharel et al., 2023) and O-GlcNAcPRED-DL (Hu et al., 2024), failed to obtain high metrics in independent test sets despite their complex architectures (an ensemble of transformers pre-trained for proteins in LM-OGlcNAc-Site’s case, and a CNN+LSTM model in O-GlcNAcPRED-DL’s case) and the direct use of some structural information. These two models still cannot reach precision $\geq$ 9% on independent test sets, and an analysis by Seber & Braatz (2024) (Table 4 of that work) noted that their model had a 2-fold performance improvement over O-GlcNAcPRED-DL, despite the fact that the latter was operating with a few advantages.
While the work of Seber & Braatz (2024) was a significant improvement over previous models and brought interpretability for the first time in a publication on data-driven O-GlcNAcylation site prediction models, the test-set F1 score and MCC achieved by the best RNN of Seber & Braatz (2024) was 36.17% and 34.57%. Thus, there is room for improvement, and we believed the transformer architecture could lead to a better model, but this hypothesis proved false, as LSTMs outperform transformers. This work also develops and explores using a novel loss function that aims to directly improve the models’ MCCs, and compares the superior results attained with new loss function against results obtained with the weighted cross-entropy loss function. 5-fold cross-validation is employed in the model training, ensuring the reliability of reported metrics and predictions. The improved model generated by this study is provided as open-source software, allowing the reproducibility of the work, the retraining of the models as additional or higher quality O-GlcNAcylation data become available, and the use of the models to further improve the understanding and applications of O-GlcNAcylation.
2 Materials and Methods
2.1 Dataset
The data used to train the model was first generated by Wulff-Fuentes et al. (2021) and was modified by Seber & Braatz (2024). The modified version is available at github.com/PedroSeber/O-GlcNAcylation_Prediction/blob/master/OVSlab_allSpecies_O-GlcNAcome_PS.csv. This dataset contains 558,168 unique S/T sites from mammalian proteins. Out of these, 13,637 (2.44%) are O-GlcNAcylated. The same procedure as in Seber & Braatz (2024) was employed to filter homologous and isoform proteins, a sequence selection based on a window size of 5 AA on each side of the central S/T (11 AA total), even for the larger windows. 20% of each dataset is separated for testing, with the remaining 80% used for cross-validation with five folds. This procedure follows what was previously done in Seber & Braatz (2024). To ensure that the choice of train/cross-validation and test sets is not biased and to highlight the robustness of our architecture and novel loss function, this work also performs nested validation with the best model, similarly to Seber & Braatz (2024). In each nested validation round, 20% of the data are used as a test set and the remaining is used for hyperparameter selection via 5-fold cross-validation (as above).
2.2 Transformer Models
[…]
2.3 RNN Models
RNN models (specifically, LSTMs) are constructed using the same packages mentioned in Section 2.2. RNN hidden sizes === {225 to 1575 in multiples of 75} for the first cell, === {0 to 300 in multiples of 75} for the second cell, and === {0 to 225 in multiples of 75} for the third and fourth cells, learning rates === {1×$\times 10^{-3}$, 5×$\times 10^{-4}$, 1×$\times 10^{-4}$}, post-LSTM MLP sizes === {37 and 75 to 825 in multiples of 75}, and window sizes of {20} were used. Training is done in 70 epochs with the AdamW optimizer with a weight decay parameter $\lambda=$0, 1×$\times 10^{-2}$, or $\times 10^{-2}$}, Kaiming uniform initialization (He et al., 2015) with $a=\sqrt{5}$, and cosine scheduling (Loshchilov & Hutter, 2017). Hyperparameter selection and testing / evaluation procedures are as in Section 2.2.
2.4 Loss Functions
The cross-entropy (CE) loss is the most common loss function for classification tasks, and its weighted variant is suitable for unbalanced data. Despite its widespread use and efficacy, the CE loss has some important limitations. First, the CE loss is distinct from the evaluation metrics used. Users of classification models may desire to maximize the F1 score or MCC metrics, but the model with the lowest CE loss is not necessarily the one with the highest metrics. Moreover, a model can lower its CE loss by manipulating the predictions of samples that are already correctly separated. For an extreme example, consider a model that predicts 0.8 for all positive samples and 0.2 for all negative samples and a second model that predicts 0.99 for all positive samples and 0.01 for all negative samples. Both models perform perfect classification, yet the second model will have a lower CE loss. In a less extreme scenario, a model may improve its CE loss by simply manipulating the probabilities of easy-to-classify samples, but such manipulations would not improve real metrics such as the F1 score or MCC.
There are multiple ways to remedy these issues. Lin et al. (2020) created the Focal Loss, which provides a correction factor $\gamma$ that reduces the effect of increasing the certainty of a correct prediction, forcing the model to improve on difficult-to-classify samples to lower the focal loss. Berman et al. (2018) created the Lovász loss, which is an optimizable form of the Jaccardi Index and is claimed to perform better than the CE loss due to its better categorizing of small objects and lower false negatives. Other losses not evaluated in this work include the Baikal loss of Gonzalez & Miikkulainen (2020), the soft cross entropy of Ilievski & Feng (2017), the sigmoid F1 of Bénédict et al. (2022), the work of Lee et al. (2021), and the MCC loss of Abhishek & Hamarneh (2021). In particular, the approaches of this work resemble those of Lee et al. (2021) and Abhishek & Hamarneh (2021), but these references did not use weighting or focal transforms, and the definitions for the differentiable F1 and MCC loss functions used in these works are slightly different from the definitions in this work.
The standard formulations of evaluation metrics (such as the F1 score or MCC) are not differentiable, so they cannot be directly used as loss functions. However, it is possible to modify them for differentiability. One method is treating the true / false categories as prediction probabilities instead of binary values. For example, if the model predicts $p_{t}=0.7$ for a positive (negative) sample, that sample would increase the true positives (negatives) by 0.7 and the false negatives (positives) by 0.3. This modification can be considered as a differentiable MCC loss function. The focal modification of Lin et al. (2020) can be added to this loss function by modifying $p_{t}$ to $(p_{t})^{\gamma}$, which leads to the focal differentiable MCC loss function. Finally, it is possible to add class weighting: by multiplying the true positives and false negatives by a scalar W, it is possible to prioritize predictions on the positive class (W > 1) or negative class (W < 1). All the above changes combined lead to the weighted focal differentiable MCC loss function (Algorithm 1).
Algorithm 1 weighted focal differentiable MCC
  Input: predictions $pred$ and targets $y$; hyperparameters $W$ (weight) and $\gamma$ (focal loss factor)
  Define the following variables: # The asterisk (*) represents multiplication. .sum() is as per the NumPy/PyTorch function. # $p_{t}$ is the probability of a correct prediction, independently of whether the true target is positive or negative.
  $p_{t}=pred*y+(1-pred)*(1-y)$
  TP = $((p_{t})^{\gamma}*y).sum()*W$
  FN = $((1-p_{t})^{\gamma}*y).sum()*W$ $W$ is used only if y = 1, that is, if the sample is in the positive class.
  FP = $((1-p_{t})^{\gamma}*(1-y)).sum()$
  TN = $((p_{t})^{\gamma}*(1-y)).sum()$
  Calculate the MCC as normal
  return -MCC # Negative for maximization
For the weighted CE loss, the weights for the positive class used were === {15, 20, 25, 30}. For the focal loss, the $\alpha$ values used were {0.9375, 0.95, 0.99, 0.999, 0.9999, 0.99999, 0.999999} and the $\gamma$ values used were {0, 1, 2, 3, 5, 10, 15, 20, 30, 45, 50}. For the weighted focal differentiable MCC, the weights W used were === {1, 2, 3, 4, 5, 10, 15, 20} and the $\gamma$ values used were {1, 2, 3, 3.5, 4, 4.5, 5}.
3 Results
3.1 Transformer Encoders Have Good Predictive Power on This Dataset, but are Inferior to RNNs
The primary goal of this subsection is to compare transformer models with the best RNN from Seber & Braatz (2024) using the modified dataset of Wulff-Fuentes et al. (2021) (Section 2.1) and the weighted CE loss function. While our transformer models surpass models published before Seber & Braatz (2024) and achieve an F1 score equal to 24.31% and an MCC equal to 22.49% (Table 1), transformers exhibit inferior recall at the same precision level (Fig. 1) than the RNNs from Seber & Braatz (2024), leading to inferior F1 and MCC metrics. It should be noted that the work of Pokharel et al. (2023) used an ensemble of three transformer models to predict O-GlcNAcylation and obtained considerably inferior F1 and MCC metrics than the RNN of Seber & Braatz (2024) and the transformer models in this work. It is possible that transformers are an ill-suited architecture for this task, potentially due to the amount of data available, which is why we train primarily RNNs in the following sections.
Figure 1: P-R curves for O-GlcNAcylation site prediction models tested on the modified dataset of Wulff-Fuentes et al. (2021). The black dotted lines are isolines of the F1 score, as labeled on the top and right sides of the plot. Model “YinOYang” is from Gupta & Brunak (2002); its metrics were obtained by Seber & Braatz (2024). Models “RNN-#” come from Seber & Braatz (2024). Models “Transformer-15” come from this work. The number after each RNN model is the LSTM module size used; the number after each transformer model is the hidden size used; the number after the semicolon for both types of models represents the window size on each side of the central S/T. The numbers on the “RNN-#” and “Transformer-15” curves represent the minimum threshold for a site to be considered positive. “No Model” sets all sites to potentially positive – that is, sets the threshold to 0.
Table 1: Performance metrics (in %) for the highest F1 score point of models tested on the modified dataset of Wulff-Fuentes et al. (2021). Models “Transformer-15” come from this work. Model “RNN-225” comes from Seber & Braatz (2024). The number after the RNN model is the LSTM module size used; the number after each transformer model is the hidden size used; the number after the semicolon for both types of models represents the window size on each side of the central S/T. Model “YinOYang” is from Gupta & Brunak (2002); its metrics were obtained by Seber & Braatz (2024). “No Model” sets all sites to potentially positive – that is, sets the threshold to 0.
Metric Transformer-15; 5 win (This Work) Transformer-15; 10 win (This Work) Transformer-15; 15 win (This Work) Best Threshold 0.60 0.20 0.50 Recall (%) 20.54 26.06 26.02 Precision (%) 15.63 13.77 21.28 F1 Score (%) 17.75 18.02 23.41 MCC (%) 15.50 16.08 21.36
Metric Transformer-15; 20 win (This Work) RNN-225; 20 win (Seber & Braatz (2024)) YinOYang (Gupta & Brunak (2002)) No Model Best Threshold 0.50 0.60 0.60 N/A Recall (%) 30.89 35.47 13.59 100 Precision (%) 20.04 36.90 8.08 2.44 F1 Score (%) 24.31 36.17 10.13 4.76 MCC (%) 22.49 34.57 7.57 0.00
A surprising result is how the performance of the transformer models does not change significantly with the hyperparameters. Using large hidden sizes or adding another linear layer after the encoders slightly degrades the performance, while increasing the number of attention heads (to 4–5) and the number of stacked encoder layers (to 4) marginally improves the model. The most surprising result is the lack of significant improvement with increasing window size. As in Seber & Braatz (2024), our models perform better with larger window sizes for sizes up to 20 (Fig. 1 and Table 1). The performance of a transformer encoder with a window size = 5 is only marginally inferior to the RNN from Seber & Braatz (2024) using the same window size. However, whereas the RNNs’ performances increase quite significantly with increasing window size, the performance of the transformers barely improves. The only significant performance increase comes from increasing the window size from 10 to 15, but barely any performance is gained from further increasing this window size to 20 (Fig. 1).
3.2 The weighted focal differentiable MCC loss function is a superior alternative to the weighted CE
We hypothesized that changing the loss function can lead to improvements in the models’ metrics, as the weighted CE is optimizing a function that is only correlated with F1 scores and MCC. Four other loss functions (as described in Section 2.4) are also used to cross-validate transformer models through the same procedure used with the weighted CE function (Sections 2.2 and 3.1).
The Focal, Lovász, and differentiable F1 losses fail to achieve meaningful results. In the best-case scenarios, models trained with these loss functions can achieve F1 scores equal to or slightly above 4.74%, equivalent to treating all (or almost all) sequences as positive. The differentiable F1, in particular, frequently gets stuck at this local F1 maximum. Naturally, these models have an MCC $\approx 0$, as they are no better than random classifiers.
Conversely, the weighted focal differentiable MCC loss is able to attain very good performance on this problem (Fig. 2 and Table 2). For both architectures, its performance is superior to that of the weighted cross-entropy loss. As in Section 3.1, RNNs are considerably better than transformer encoders in predicting O-GlcNAcylation. The RNN model trained with the weighted focal differentiable MCC loss achieves an F1 score = 37.03% and an MCC = 36.58%. The relative and absolute increases in MCC are higher than those for the F1 score, highlighting the effectiveness of the weighted focal differentiable MCC loss, as it seeks to directly optimize the MCC. The weighted focal differentiable MCC loss can also be used for fine tuning; an RNN model first trained using weighted cross-entropy, then fine-tuned using the weighted focal differentiable MCC is generated (“RNN-225; CE $\rightarrow$ diff MCC”). This model initially followed the optimal hyperparameters (including model size) found through cross-validation using the weighted cross-entropy loss, but the learning rate and loss hyperparameters are changed to the optimal values found through cross-validation with the weighted focal differentiable MCC loss function when that loss function is used for fine-tuning. This fine-tuning also leads to a model that is better than the previous state-of-the-art from Seber & Braatz (2024) (Fig. 2 and Table 2), reaching an F1 score = 36.52% and an MCC = 36.01%, values slightly lower than that of the model trained only with the weighted focal differentiable MCC loss.
There is a significant difference between the optimal hyperparameters selected through the weighted cross-entropy and weighted focal differentiable MCC loss functions. Compared to the optimal values found through the weighted cross-entropy loss, the weighted focal differentiable MCC loss has an optimum at higher model hidden sizes (15 vs. 240 for transformers; 225 vs. 1425 for RNNs), higher batch sizes (256 vs. 512 for transformers; 32 vs. 128 for RNNs), lower learning rates (1×$\times 10^{-3}$ vs. 1×$\times 10^{-4}$), and lower class weights (15 vs. 3 for transformers; 30 vs. 3 for RNNs).
Figure 2: P-R curves for O-GlcNAcylation site prediction models tested on the modified dataset of Wulff-Fuentes et al. (2021). The black dotted lines are isolines of the F1 score, as labeled on the top and right sides of the plot. Model “RNN-225; 20 win; CE” comes from Seber & Braatz (2024). All others come from this work. The number after each RNN model is the LSTM module size used; the number after each transformer model is the hidden size used; the number after the semicolon for both types of models represents the window size on each side of the central S/T. The numbers on the curves represent the minimum threshold for a site to be considered positive. “No Model” sets all sites to potentially positive – that is, sets the threshold to 0.
Table 2: Recall, precision, F1 score, and MCC metrics (in %) for the highest F1 score point of models tested on the modified dataset of Wulff-Fuentes et al. (2021). Model “RNN-225; CE” comes from Seber & Braatz (2024). All others come from this work. The number after each RNN model is the LSTM module size used; the number after each transformer model is the hidden size used; the number after the semicolon for both types of models represents the window size on each side of the central S/T. “CE” represents the weighted cross-entropy loss, and “diff MCC” represents the weighted focal differentiable MCC loss. “No Model” sets all sites to potentially positive – that is, sets the threshold to 0.
Metric Transformer-15; CE (This Work) Transformer-240; diff MCC (This Work) RNN-225; CE (Seber & Braatz (2024)) Best Threshold 0.50 0.40 0.60 Recall (%) 30.89 25.70 35.47 Precision (%) 20.04 25.48 36.90 F1 Score (%) 24.31 25.59 36.17 MCC (%) 22.49 23.67 34.57
Metric RNN-1425; diff MCC (This Work) RNN-225; CE $\rightarrow$ diff MCC (This Work) No Model Best Threshold 0.70 0.40 N/A Recall (%) 30.67 30.35 100 Precision (%) 46.70 45.84 2.44 F1 Score (%) 37.03 36.52 4.76 MCC (%) 36.58 36.01 0.00
3.3 Stacked RNNs provide better performance, and the weighted focal differentiable MCC loss is still superior
Seber & Braatz (2024) trained RNNs with only a single cell, a restriction that was followed by this work in Section 3.2. However, RNNs with multiple stacked cells may perform better than single-cell RNNs. To verify this hypothesis, RNN models with 2, 3, and 4 LSTM layers are tested (as per Section 2.3). Similarly to Section 3.2, three models are created: one trained only with the weighted CE loss, one trained only with the weighted focal differentiable MCC loss, and one first trained with the weighted CE loss, then fine-tuned with the weighted focal differentiable MCC loss. These models all surpass their single-RNN layer equivalents from Section 3.2. The model trained with the weighted CE loss (“RNN-[450,75]; CE”) displays the greatest relative improvement over its single-cell version, reaching an F1 score = 38.73 % and an MCC = 37.31%. Like with the single-cell RNNs, fine-tuning with the weighted focal differentiable MCC loss after training with the weighted CE leads to improvements in the model performance; this fine-tuned model (“RNN-[450,75]; CE $\rightarrow$ diff MCC”) reaches an F1 score = 38.83% and an MCC = 37.33%. While the performance improvement of fine-tuning is smaller in this scenario, these improved results again confirm the efficacy of fine-tuning with the weighted focal differentiable MCC loss. Finally, a model trained directly with the weighted focal differentiable MCC loss (“RNN-[600,75]; diff MCC”) achieves the best results, reaching an F1 score = 38.88% and an MCC = 38.20%. The increase in MCC is considerably higher, further corroborating the benefits of this new loss function discussed in Section 3.2. Overall, the “RNN-[600,75]; diff MCC” model displays a 7.5% increase in F1 score and a 10.5% increase in MCC over the “RNN-255; CE” model of Seber & Braatz (2024).
Figure 3: P-R curves for two-layer O-GlcNAcylation site prediction models tested on the modified dataset of Wulff-Fuentes et al. (2021). The black dotted lines are isolines of the F1 score, as labeled on the top and right sides of the plot. The number after each RNN model is the LSTM module size used, and the number after the semicolon represents the window size on each side of the central S/T. The numbers on the curves represent the minimum threshold for a site to be considered positive. “No Model” sets all sites to potentially positive – that is, sets the threshold to 0.
Table 3: Recall, precision, F1 score, and MCC metrics (in %) for the highest F1 score point of two-layer models tested on the modified dataset of Wulff-Fuentes et al. (2021). The number after each RNN model is the LSTM module size used, and the number after the semicolon represents the window size on each side of the central S/T. “CE” represents the weighted cross-entropy loss, and “diff MCC” represents the weighted focal differentiable MCC loss. The “No Model” (threshold = 0) baseline metrics are as in Tables 1 and 2.
Metric RNN-[450,75]; CE (This Work) RNN-[450,75]; CE $\rightarrow$ diff MCC (This Work) RNN-[600,75]; diff MCC (This Work) Best Threshold 0.90 0.50 0.60 Recall (%) 36.69 37.62 33.11 Precision (%) 41.02 40.11 47.10 F1 Score (%) 38.73 38.83 38.88 MCC (%) 37.31 37.33 38.20
The models with the highest average cross-validation F1 score all had only two layers and second LSTM layers with 75 neurons. As in Section 3.2, the optimal hyperparameters for models trained with the weighted focal differentiable MCC loss included higher model hidden sizes (600 vs. 450 for CE), lower learning rates (5×$\times 10^{-4}$ vs. 1×$\times 10^{-3}$), and lower class weights (2 vs. 20). Surprisingly, both models had the same optimal batch size (128), and the hidden size and class weights for the model trained with the weighted CE loss moved towards those from the model trained with the weighted focal differentiable MCC loss.
Finally, we verified that our splits for the cross-validation and test sets do not bias the models by performing five-fold nested cross-validation, following the procedure in Section 2.1, on the “RNN-[600,75]; diff MCC” model. The models generated by each fold had very similar performance to each other (Fig. A1 and Table A1), highlighting the robustness of that architecture and loss function to changes in the training/test data.
4 Discussion
This work generates transformer encoder (Section 3.1) and RNN (Sections 3.2 and 3.3) models to predict O-GlcNAcylation sites based on protein sequences using a modified version of an extensive public dataset Wulff-Fuentes et al. (2021). These models are also compared to the state-of-the-art RNNs trained by Seber & Braatz (2024).
Although transformer models attain high performances on this dataset, they are inferior to the RNN models trained by Seber & Braatz (2024), reaching an F1 score of 24.31% and an MCC of 22.49% (Section 3.1, Fig. 1, and Table 1). A significant difference between transformers and RNNs is that the former do not scale well with increasing window sizes, while the latter display significant improvements with larger window sizes. With window size = 5, both transformers and RNNs possess similar performances, but a large performance gap arises with increasing window sizes. Multiple reasons could be causing this inferior transformer performance, including a lack of training data. Although the dataset from Wulff-Fuentes et al. (2021) is one of the largest (if not the largest) O-GlcNAcylation datasets published, it is possible that there simply are not enough data for the transformer architecture to perform well in this task.
Because the weighted cross-entropy loss function, which is used to train the models in Seber & Braatz (2024) and Section 3.1 of this work, only indirectly optimizes the classification performance metrics, we hypothesized that a function that could directly improve the F1 score and MCC can lead to better-performing models.
We develop the weighted focal differentiable MCC loss to address this question (Section 2.4; Algorithm 1). This function allows the direct optimization of the MCC, yielding better classification models. We cross-validate and test transformer and RNN models using the same dataset and notice that the performance of the weighted focal differentiable MCC loss is similar to but higher than that of the weighted cross-entropy loss (Section 3.2, Fig. 2, and Table 1). This loss function improves both the transformer and RNN architectures, and leads to the creation of an improved model with an F1 score = 37.03% and an MCC = 36.58%. Furthermore, a model first trained using the weighted cross-entropy loss, then fine-tuned using the weighted focal differentiable MCC loss achieves improved performance in this prediction task relative to a model trained using only the weighted CE loss, reaching an F1 score = 36.52% and an MCC = 36.01%. While these metrics are slightly smaller than the values obtained by a model directly trained with the weighted focal differentiable MCC loss, these results show how this new loss function can also be used to optimize and improve trained models at a low cost. The optimal hyperparameters for models trained with the weighted focal differentiable MCC loss are different from those of models trained with the weighted cross-entropy loss.
As Seber & Braatz (2024) used only single-cell models in their work, the RNN models from Section 3.2 follow this restriction. Models with multiple stacked cells can have better performance, so we train RNN models with multiple LSTM cells in Section 3.3. These models display superior cross-validation F1 scores over models with one cell, and the best models have only 75 neurons in their 2nd cells. No matter the loss function used, these deeper models display better performance in the test set than the best RNN of Seber & Braatz (2024) (Fig. 3 and Table 3). As in Section 3.2, fine-tuning using the weighted focal differentiable MCC loss after training with the weighted CE leads to improvements in performance over using only the weighted CE, corroborating one of the benefits of the weighted focal differentiable MCC loss. The model trained solely with the weighted focal differentiable MCC loss, “RNN-[600,75]; diff MCC”, achieves state-of-the-art performance in O-GlcNAcylation prediction, reaching an F1 score = 38.88% and an MCC = 38.20%. Overall, this is a 7.5% increase in F1 score and a 10.5% increase in MCC over the “RNN-255; CE” model of Seber & Braatz (2024).
There are two main limitations to this work. On the O-GlcNAcylation prediction side, although the models developed in this work reach state-of-the-art performance in multiple independent datasets, that performance is still far away from (near-)perfect, possibly limiting the widespread usability of the models. It should be noted that the experimental detection of O-GlcNAcylation sites is a laborious and expensive process, so even an imperfect model may be useful. Having a usable model is particularly important given the relevance of O-GlcNAcylation to many diseases and the potential of O-GlcNAcylation as a target for medicines (Section 1). On the weighted focal differential MCC loss side, although we earnestly believe that it can improve the performance of models in almost any classification task relative to the (weighted) CE, this work tests its performance on only one task. Future directions primarily include testing the weighted focal differential MCC loss on multiple classification tasks to show its wide usability, including in fine-tuning models that were previously trained with the CE loss. Another direction is analyzing the effects of dynamically changing this loss function’s hyperparameters over the epochs to determine whether it can be further improved.
The code used in this work is publicly available, allowing the reproduction, improvement, and reuse of this work. It is simple to install and run the best RNN model to predict O-GlcNAcylation sites based on the local protein sequence. Instructions are provided in Sections A1.1 and A1.2 of the supplemental data or the README in our GitHub <redacted for review> repository.

### v3

@@
 Abstract
-Glycosylation, a protein modification, has multiple essential functional and structural roles.
 O-GlcNAcylation, a subtype of glycosylation, has the potential to be an important target for therapeutics, but methods to reliably predict O-GlcNAcylation sites had not been available until 2023; a 2021 review correctly noted that published models were insufficient and failed to generalize.
 Moreover, many are no longer usable.
-In 2023, a considerably better recurrent neural network (RNN) model with an F1 score of 36.17% and a Matthews Correlation Coefficient (MCC) of 34.57% on a large dataset was published.
-This article first sought to improve these metrics using transformer encoders.
-While transformers displayed high performance on this dataset, their performance was inferior to that of the previously published RNN.
-We then created a new loss function, which we call the weighted focal differentiable MCC, to improve the performance of classification models.
-RNN models trained with this new function display superior performance to models trained using the weighted cross-entropy loss; this new function can also be used to fine-tune trained models.
-A two-cell RNN trained with this loss achieves state-of-the-art performance in O-GlcNAcylation site prediction with an F1 score of 38.88% and an MCC of 38.20% on that same large dataset.
-1 Introduction
-Glycosylation, a co- and post-translational modification, occurs when glycan(s) are added to proteins.
-O-linked glycosylation occurs when a glycan is added to an oxygen of an amino acid (usually serine or threonine in mammals).
-O-GlcNAcylation, a subtype of O-linked glycosylation, occurs when the first glycan added is an N-Acetylglucosamine (abbreviated GlcNAc) (Schjoldager et al., 2020).
+In 2023, a considerably better recurrent neural network (RNN) model was published.
+This article creates improved models by using a new loss function, which we call the weighted focal differentiable MCC.
+RNN models trained with this new loss display superior performance to models trained using the weighted cross-entropy loss; this new function can also be used to fine-tune trained models.
+An RNN trained with this loss achieves state-of-the-art performance in O-GlcNAcylation site prediction with an F1 score of 38.88% and an MCC of 38.20% on an independent test set from the largest dataset available.
+Author Summary
+Glycosylation is the attachment of a sugar or sugars to a protein.
+This natural process has many essential functions in organisms.
+O-GlcNAcylation is a subtype of glycosylation that has been identified as having high potential for therapeutics.
+However, no method for systematic identification of O-GlcNAcylation sites exists.
+Models have been used for this task, but a 2021 review correctly noted that many models have low performances when predicting on data outside their training sets, making them unsuitable for practical use.
+A 2023 work made a considerable improvement in models for O-GlcNAcylation, leading to the first usable model for O-GlcNAcylation site prediction.
+We create improved models, which have a greater ability to identify O-GlcNAcylation sites, in turn better assisting the development of related therapeuticals and increasing our understanding of the biology behind this process.
+We also create a new loss function for classification models.
+Machine-learning models are optimized by minimizing a loss function, which represents the difference between the model’s current performance and a perfect performance.
+The most widely used loss function only indirectly optmizes a model’s performance metrics, but our loss function can directly optmize important metrics, leading to better performance, including for models made for other (non-glycosylation) tasks.
+Introduction
+Glycosylation, a co- and post-translational modification, occurs when glycans are added to proteins.
+O-linked glycosylation occurs when a glycan is added to an oxygen of an amino acid (usually serine or threonine in mammalian proteins).
+O-GlcNAcylation, a subtype of O-linked glycosylation, occurs when the first glycan added is an N-Acetylglucosamine (abbreviated GlcNAc) [1].
 Unlike other forms of glycosylation, O-GlcNAcylation may be viewed similarly to phosphorylation, for no other glycans are further added to a site after GlcNAc and the process is catalyzed solely by two enzymes, OGT and OGA.
-O-GlcNAcylation is important functionally and structurally (Schjoldager et al., 2020; Chang et al., 2020); conversely, incorrect O-GlcNAcylation or its improper removal is associated with multiple diseases such as cancers (Shi et al., 2022), infections (Chang et al., 2020), and heart failure (Umapathi et al., 2021).
-According to recent research, O-GlcNAcylation can be a powerful target for therapeutics (Zhu & Hart, 2021), further emphasizing its relevance.
-While O-GlcNAcylation is essential for human health and very important for biotherapeutics, challenges still exist.
+O-GlcNAcylation is important functionally and structurally [1, 2]; conversely, incorrect O-GlcNAcylation or its improper removal is associated with multiple diseases such as cancers [3], infections [2], and heart failure [4].
+According to recent research, O-GlcNAcylation can be a powerful target for therapeutics [5], further emphasizing its relevance.
+While O-GlcNAcylation is essential for human health and can be critical for biotherapeutics, challenges still exist.
 Because O-GlcNAcylation is complete after the addition of a single glycan, the problem of predicting glycan distributions in a glycosylation site does not exist; thus, the main predictive task is to determine where and when (if ever) an amino acid will be O-GlcNAcylated.
-As summarized by Seber & Braatz (2024), this task is challenging for “multiple reasons”, which include “a low frequency of events (only about 2% of S/T sites are O-GlcNAcylated) and a lack of a motif to guide predictive efforts”.
-Moreover, Seber & Braatz (2024) also note that “the effects of neighboring amino acids likely influence whether an S/T is O-GlcNAcylated”, a fact corroborated by their interpretability studies done with their final model, but many previous works use model architectures that cannot directly take into account any type of sequential information.
-The pioneer classifier machine-learning model for this challenging task is YinOYang (YoY) (Gupta & Brunak, 2002), which also had one of the highest performances until the work of Seber & Braatz (2024), first published in 2023.
-However, until at least 2021, O-GlcNAcylation site prediction models had insufficient performance to help advance research in the area (Mauri et al., 2021).
-Specifically, Mauri et al.
-(2021) found that no published model until then could achieve a precision $\geq$ 9% on a medium-sized independent dataset, suggesting that models to predict the location of O-GlcNAcylation sites failed to generalize successfully in spite of the high metrics achieved with their training data.
-The models evaluated by Mauri et al.
-(2021) also have low F1 scores and Matthew Correlation Coefficients (MCCs), another sign that their performance is insufficient.
-Seber & Braatz (2024) used recurrent neural network (RNN) models (specifically, long short-term memory [LSTM] models) to predict the presence of O-GlcNAc sites from mammalian protein sequence data.
-That work used two different datasets (from Mauri et al.
-(2021); Wulff-Fuentes et al.
-(2021)), the latter of which is among the largest O-GlcNAcylation datasets available.
+As summarized by Ref.
+[6], this task is challenging for “multiple reasons”, which include “a low frequency of events (only about 2% of S/T [serine/threonine] sites are O-GlcNAcylated) and a lack of a motif to guide predictive efforts”.
+Moreover, Ref.
+[6] also note that “the effects of neighboring amino acids likely influence whether an S/T is O-GlcNAcylated”, a fact corroborated by their interpretability studies done with their final model, but many previous works use model architectures that cannot directly take into account any type of sequential information.
+The pioneer classifier machine-learning model for this challenging task is YinOYang (YoY) [7], which also had one of the highest performances until the work of Ref.
+[6], first published in 2023.
+However, until at least 2021, O-GlcNAcylation site prediction models had insufficient performance to help advance research in the area [8].
+Specifically, Ref.
+[8] found that no published model until then could achieve a precision $\geq$ 9% on a medium-sized independent dataset, suggesting that models to predict the location of O-GlcNAcylation sites failed to generalize successfully in spite of the high metrics achieved with their training data.
+The models evaluated by Ref.
+[8] also have low F1 scores and Matthews Correlation Coefficients (MCCs), another sign that their performance is insufficient.
+Ref.
+[6] used recurrent neural network (RNN) models (specifically, long short-term memory [LSTM] models) to predict the presence of O-GlcNAc sites from mammalian protein sequence data.
+That work used two different datasets (from Refs.
+[8, 9]), the latter of which is among the largest O-GlcNAcylation datasets available.
 These RNNs obtained vastly superior performance to the previously published models, achieving an F1 score more than 3.5-fold higher and an MCC more than 4.5-fold higher than the previous state of the art.
-Furthermore, Shapley values were used to interpret the predictions of that RNN model through the sum of simple linear coefficients (Shapley, 1951).
-These predictions with Shapley values maintained most of the performance of the original RNN models (Supplemental data of Seber & Braatz (2024)).
-Conversely, two other models that were also released recently, LM-OGlcNAc-Site (Pokharel et al., 2023) and O-GlcNAcPRED-DL (Hu et al., 2024), failed to obtain high metrics in independent test sets despite their complex architectures (an ensemble of transformers pre-trained for proteins in LM-OGlcNAc-Site’s case, and a CNN+LSTM model in O-GlcNAcPRED-DL’s case) and the direct use of some structural information.
-These two models still cannot reach precision $\geq$ 9% on independent test sets, and an analysis by Seber & Braatz (2024) (Table 4 of that work) noted that their model had a 2-fold performance improvement over O-GlcNAcPRED-DL, despite the fact that the latter was operating with a few advantages.
-While the work of Seber & Braatz (2024) was a significant improvement over previous models and brought interpretability for the first time in a publication on data-driven O-GlcNAcylation site prediction models, the test-set F1 score and MCC achieved by the best RNN of Seber & Braatz (2024) was 36.17% and 34.57%.
-Thus, there is room for improvement, and we believed the transformer architecture could lead to a better model, but this hypothesis proved false, as LSTMs outperform transformers.
-This work also develops and explores using a novel loss function that aims to directly improve the models’ MCCs, and compares the superior results attained with new loss function against results obtained with the weighted cross-entropy loss function.
+Furthermore, Shapley values were used to interpret the predictions of that RNN model through the sum of simple linear coefficients [10].
+These predictions with Shapley values maintained most of the performance of the original RNN models (Supplemental data of Ref.
+[6]) and provided interpretability for the first time in a publication on data-driven O-GlcNAcylation site prediction models.
+Conversely, two other models that were also released recently, LM-OGlcNAc-Site [11] and O-GlcNAcPRED-DL [12], failed to obtain high metrics in independent test sets despite their complex architectures (an ensemble of transformers pre-trained for proteins in LM-OGlcNAc-Site’s case, and a CNN+LSTM model in O-GlcNAcPRED-DL’s case) and the direct use of some structural information.
+These two models still cannot reach precision $\geq$ 9% on independent test sets, and an analysis by Ref.
+[6] (Table 4 of that work) noted that their model had a 2-fold performance improvement over O-GlcNAcPRED-DL, despite the fact that the latter was operating with a few advantages.
+This work creates improved O-GlcNAcylation prediction models by using three methods: transformer encoder models directly trained on O-GlcNAcylation data, finetuning of a pre-trained protein transformer [13], both of which had lower performances than LSTM models, and transformers and RNNs trained with a new loss function, which we call the weighted focal differentiable MCC.
+This novel loss function aims to directly improve the models’ MCCs, attaining superior results relative to those obtained with the weighted cross-entropy loss function.
+LSTMs trained with this new loss function reach state-of-the-art O-GlcNAcylation site prediction results, and the improvement in MCC compared to the cross-entropy loss function is statistically significant.
 5-fold cross-validation is employed in the model training, ensuring the reliability of reported metrics and predictions.
-The improved model generated by this study is provided as open-source software, allowing the reproducibility of the work, the retraining of the models as additional or higher quality O-GlcNAcylation data become available, and the use of the models to further improve the understanding and applications of O-GlcNAcylation.
-2 Materials and Methods
-2.1 Dataset
-The data used to train the model was first generated by Wulff-Fuentes et al.
-(2021) and was modified by Seber & Braatz (2024).
-The modified version is available at github.com/PedroSeber/O-GlcNAcylation_Prediction/blob/master/OVSlab_allSpecies_O-GlcNAcome_PS.csv.
-This dataset contains 558,168 unique S/T sites from mammalian proteins.
+The improved model and novel loss function generated by this study are provided as open-source software, allowing the reproducibility of the work, the retraining of the models as additional or higher quality O-GlcNAcylation data become available, the use of the models to further improve the understanding and applications of O-GlcNAcylation, and the use of this novel loss function to improve models in other relevant tasks, as already done in Ref.
+[14], for example.
+Results
+Transformers Have Good Predictive Power, but Are Inferior to RNNs
+First, transformer encoder models were trained directly on this task.
+Although these models surpassed multiple models previously reported in the literature, they were inferior to the RNN of Ref.
+[6] (Section S4.1).
+Next, the ProteinBERT transformer was finetuned on this task.
+These finetuned transformers performed better than the transformer encoders directly trained on the O-GlcNAcylation data, but slightly worse than the RNN of Ref.
+[6] (Table 1 and Fig.
+1).
+Nevertheless, these finetuned transformers achieve good performance, surpassing multiple other models in the literature, including other transformer-based models (such as the previously mentioned LM-OGlcNAc-Site [11], which has a 2.3-fold lower MCC on an independent test set).
+The performance of finetuned ProteinBERT models would be state-of-the-art if not for the RNN of Ref.
+[6] and the RNNs trained in this work.
+Surprisingly, the finetuned ProteinBERT models’ performances began to saturate and be maximized with a window size of 15, whereas other model types had this saturation and maximization only with a window size of 20.
+As with the other models, a window size of 25 or 30 did not improve performance in a statistically significant way relative to a window size of 20.
+It is possible that transformers are an ill-suited architecture for this task, potentially due to the amount of data available, which is why we train primarily RNNs in the following sections.
+The work of Ref.
+[13], which trained ProteinBERT, also noted that LSTMs surpassed transformer models in multiple non-glycosylation-related tasks of biological relevance, corroborating this suitability hypothesis.
+Fig 1: P-R curves for the first seed of O-GlcNAcylation site prediction finetuned ProteinBERT models tested on the modified dataset of Ref.
+[9].
+The black dotted lines are isolines of the F1 score, as labeled on the top and right sides of the plot.
+The number after the semicolon represents the window size on each side of the central S/T.
+The numbers on the curves represent the minimum threshold for a site to be considered positive.
+The “RNN-225” model [6] is the previous state-of-the-art model and is included for comparison.
+“No Model” is the baseline (threshold = 0).
+Table 1: Performance metrics (in %) for the highest F1 score point of finetuned ProteinBERT models tested on the modified dataset of Ref.
+[9].
+WS is the window size on each side of the central S/T.
+The “RNN-225” model [6] is the previous state-of-the-art model and is included for comparison.
+Baseline (threshold = 0) metrics are precision = 2.44%, F1 score = 4.76%, and MCC = 0%.
+[…]
+RNNs + Our Loss Function Have State-of-the-Art Predictive Performance
+With the weighted focal differentiable MCC loss, we then trained an RNN model and also finetuned the RNN model from Ref.
+[6].
+Ref.
+[6] trained only single-cell RNNs, and so this constraint has been maintained for the first set of RNNs trained in this work.
+The RNN trained directly with our novel loss function (“RNN-1425; diff MCC”) achieves an F1 score = 37.03% and an MCC = 36.58% (Table 2 and Fig.
+2).
+The relative and absolute increases in MCC are higher than those for the F1 score, highlighting the effectiveness of the weighted focal differentiable MCC loss, as it directly optimizes the MCC.
+To show that the weighted focal differentiable MCC loss can also be used for finetuning, the RNN model from Ref.
+[6], which was first trained using weighted cross-entropy, was finetuned using the weighted focal differentiable MCC loss (“RNN-225; CE $\rightarrow$ diff MCC”).
+This model initially followed the optimal hyperparameters (including model size) found through cross-validation using the weighted cross-entropy loss, but the learning rate and loss hyperparameters are changed to the optimal values found through cross-validation with the weighted focal differentiable MCC loss function when that loss function is used for fine-tuning.
+This fine-tuning also leads to a model that is better than the previous state-of-the-art from Ref.
+[6] (Fig.
+2 and Table 2), reaching an F1 score = 36.52% and an MCC = 36.01%, values slightly lower than that of the model trained only with the weighted focal differentiable MCC loss.
+Fig 2: P-R curves for O-GlcNAcylation site prediction single-cell RNN models tested on the modified dataset of Ref.
+[9].
+Labels, curve descriptions, and baseline metrics are as in the captions of Table 1 and Fig.
+1.
+Table 2: Performance metrics (in %) for the highest F1 score point of single-cell RNN models tested on the modified dataset of Ref. [9]. Labels and baseline metrics are as in the caption of Table 1.
+Metric RNN-225; CE (Ref.
+[6]) RNN-1425; diff MCC (This Work) RNN-225; CE $\rightarrow$ diff MCC (This Work) Best Threshold 0.60 0.70 0.40 Recall (%) 35.47 30.67 30.35 Precision (%) 36.90 46.70 45.84 F1 Score (%) 36.17 37.03 36.52 MCC (%) 34.57 36.58 36.01
+[…]
+RNNs with multiple stacked cells may perform better than single-cell RNNs.
+To verify this hypothesis, RNN models with 2, 3, and 4 LSTM cells are tested (as per Section “Transformer and RNN Models”).
+As with the single-cell RNNs, three models are created: one trained only with the weighted CE loss, one trained only with the weighted focal differentiable MCC loss, and one first trained with the weighted CE loss, then fine-tuned with the weighted focal differentiable MCC loss.
+These models all surpass their single-cell equivalents (Table 3 and Fig.
+3).
+The model trained with the weighted CE loss (“RNN-[450,75]; CE”) displays the greatest relative improvement over its single-cell version, reaching an F1 score = 38.73 % and an MCC = 37.31%.
+Like with the single-cell RNNs, fine-tuning with the weighted focal differentiable MCC loss after training with the weighted CE leads to improvements in the model performance; this fine-tuned model (“RNN-[450,75]; CE $\rightarrow$ diff MCC”) reaches an F1 score = 38.83% and an MCC = 37.33%.
+While the performance improvement of fine-tuning is smaller in this scenario, these improved results again confirm the efficacy of fine-tuning with the weighted focal differentiable MCC loss.
+Finally, a model trained directly with the weighted focal differentiable MCC loss (“RNN-[600,75]; diff MCC”) achieves the best results, reaching an F1 score = 38.88% and an MCC = 38.20%.
+The increase in MCC is considerably higher, further corroborating the benefits of this new loss function discussed above.
+Overall, the “RNN-[600,75]; diff MCC” model displays a 7.5% increase in F1 score and a 10.5% increase in MCC over the “RNN-255; CE” model of Ref.
+[6].
+Fig 3: P-R curves for O-GlcNAcylation site prediction two-cell RNN models tested on the modified dataset of Ref.
+[9].
+Labels, curve descriptions, and baseline metrics are as in the captions of Table 1 and Fig.
+1.
+Table 3: Performance metrics (in %) for the highest F1 score point of two-cell RNN models tested on the modified dataset of Ref. [9]. Labels and baseline metrics are as in the caption of Table 1.
+Metric RNN-[450,75]; CE (This Work) RNN-[450,75]; CE $\rightarrow$ diff MCC (This Work) RNN-[600,75]; diff MCC (This Work) Best Threshold 0.90 0.50 0.60 Recall (%) 36.69 37.62 33.11 Precision (%) 41.02 40.11 47.10 F1 Score (%) 38.73 38.83 38.88 MCC (%) 37.31 37.33 38.20
+[…]
+Finally, we verified that our splits for the cross-validation and test sets do not bias the models by performing five-fold nested cross-validation, following the procedure in Section “Dataset”, on the “RNN-[600,75]; diff MCC” model.
+The models generated by each fold had very similar performance to each other (Fig.
+S2 and Table S2), highlighting the robustness of that architecture and loss function to changes in the data.
+In addition, a comparison of the nested validation results obtained with the weighted focal differentiable MCC loss function (Table S2) and those obtained with the CE loss function (Table S3) show that the increases in MCC brought by the use of the diffMCC loss function are statistically significant (p = 0.0115) over multiple test sets, corroborating the success of our novel loss function in improving models by directly optimizing their MCCs.
+Discussion
+This work generates transformer encoder (Section S4.1), finetuned transformer (Section “Transformers Have Good Predictive Power, but Are Inferior to RNNs”), and RNN (Section “RNNs + Our Loss Function Have State-of-the-Art Predictive Performance”) models to predict O-GlcNAcylation sites based on protein sequences using a modified version of an extensive public dataset [9].
+These models are also compared to the state-of-the-art RNNs trained by Ref.
+[6].
+As the weighted cross-entropy loss function only indirectly optimizes the classification performance metrics, we hypothesized that a function that could directly improve the F1 score and MCC can lead to better-performing models.
+We develop the weighted focal differentiable MCC loss to address this question (Section “Loss Functions”; Algorithm 1).
+This function directly optimizes the MCC, yielding better classification models.
+This new loss function was used to generate a state-of-the-art model for O-GlcNAcylation site prediction, reaching an F1 score = 38.88% and an MCC = 38.20% (Fig.
+3 and Table 3).
+Overall, this is a 7.5% increase in F1 score and a 10.5% increase in MCC over the “RNN-255; CE” model of Ref.
+[6].
+This improved model can assist in the discovery of novel O-GlcNAcylation sites, in turn expanding our knowledge of glycan biology and enabling the discovery of O-GlcNAcylation-related drugs [5].
+As the identification of O-GlcNAcylation sites is a laborious, specialized, and time-consuming process, the use of models and computational tools can accelerate research in the field.
+Furthermore, the creation of a new loss function for classification tasks (Algorithm 1) can enable the training of better models, as classification is ubiquitous in many scientific, engineering, and social fields.
+This loss function improved models in a statistically significant way (p = 0.0115).
+These improvements occur not just for models used for O-GlcNAcylation-related tasks, as evidenced by the successful use of our weighted focal differentiable MCC loss function in other works [14].
+The code used in this work is freely and publicly available, allowing the reproduction, improvement, and reuse of this work.
+It is simple to install and run the best RNN model to predict O-GlcNAcylation sites based on the local protein sequence.
+Instructions are provided in Sections S1 and S2 of the supplemental data or the README at https://github.com/PedroSeber/
+O-GlcNAcylation_Prediction
+Methods
+Dataset
+The data used to train the model was first generated by Ref. [9] and was modified by Ref. [6]. The modified version is available at
+https://github.com/PedroSeber/O-GlcNAcylation_Prediction/blob/master/
+OVSlab_allSpecies_O-GlcNAcome_PS.csv
+.
+This dataset contains 558,168 unique serine/threonine (S/T) sites from mammalian proteins.
 Out of these, 13,637 (2.44%) are O-GlcNAcylated.
-The same procedure as in Seber & Braatz (2024) was employed to filter homologous and isoform proteins, a sequence selection based on a window size of 5 AA on each side of the central S/T (11 AA total), even for the larger windows.
-20% of each dataset is separated for testing, with the remaining 80% used for cross-validation with five folds.
-This procedure follows what was previously done in Seber & Braatz (2024).
-To ensure that the choice of train/cross-validation and test sets is not biased and to highlight the robustness of our architecture and novel loss function, this work also performs nested validation with the best model, similarly to Seber & Braatz (2024).
+The same procedure as in Ref.
+[6] was employed to filter homologous and isoform proteins, a sequence selection based on a window size of 5 amino acids (AAs) on each side of the central S/T (11 AAs total), even for the larger windows.
+20% of each dataset is separated for testing, with the remaining 80% used for cross-validation with five folds; a procedure also previously done in Ref.
+[6].
+To ensure that the choice of train/cross-validation and test sets is not biased and to highlight the robustness of our architecture and novel loss function, this work also performs nested validation with the best model, similarly to Ref.
+[6].
 In each nested validation round, 20% of the data are used as a test set and the remaining is used for hyperparameter selection via 5-fold cross-validation (as above).
-2.2 Transformer Models
-[…]
-2.3 RNN Models
-RNN models (specifically, LSTMs) are constructed using the same packages mentioned in Section 2.2.
-RNN hidden sizes === {225 to 1575 in multiples of 75} for the first cell, === {0 to 300 in multiples of 75} for the second cell, and === {0 to 225 in multiples of 75} for the third and fourth cells, learning rates === {1×$\times 10^{-3}$, 5×$\times 10^{-4}$, 1×$\times 10^{-4}$}, post-LSTM MLP sizes === {37 and 75 to 825 in multiples of 75}, and window sizes of {20} were used.
-Training is done in 70 epochs with the AdamW optimizer with a weight decay parameter $\lambda=$0, 1×$\times 10^{-2}$, or $\times 10^{-2}$}, Kaiming uniform initialization (He et al., 2015) with $a=\sqrt{5}$, and cosine scheduling (Loshchilov & Hutter, 2017).
-Hyperparameter selection and testing / evaluation procedures are as in Section 2.2.
-2.4 Loss Functions
+Transformer and RNN Models
+Transformer encoder and RNN (specifically, LSTM) models are constructed using PyTorch [15] and other Python packages [16, 17, 18].
+Moreover, the ProteinBERT transformer model [13] was also finetuned for this task.
+Details on the hyperparameters used for each model can be found in Section S3.
+The best combination of hyperparameters is determined by a grid search, and the combination with the highest cross-validation average F1 score is selected.
+Finally, the performance on an independent test set of a model using that combination of hyperparameters is reported.
+Multiple thresholds are used to construct a Precision-Recall (P-R) curve, and the F1 scores and MCCs of the models are analyzed, allowing for a throughout evaluation of the best model’s performance.
+F1 scores and MCCs are used for evaluation because they are widely regarded as the best performance metrics for classification models, as discussed in Section 2.3 of Ref.
+[6] and references within.
+Loss Functions
 The cross-entropy (CE) loss is the most common loss function for classification tasks, and its weighted variant is suitable for unbalanced data.
@@
 Users of classification models may desire to maximize the F1 score or MCC metrics, but the model with the lowest CE loss is not necessarily the one with the highest metrics.
-Moreover, a model can lower its CE loss by manipulating the predictions of samples that are already correctly separated.
+Moreover, a model can lower its CE loss by manipulating the predictions of samples that are already correctly classified.
 For an extreme example, consider a model that predicts 0.8 for all positive samples and 0.2 for all negative samples and a second model that predicts 0.99 for all positive samples and 0.01 for all negative samples.
@@
 There are multiple ways to remedy these issues.
-Lin et al.
-(2020) created the Focal Loss, which provides a correction factor $\gamma$ that reduces the effect of increasing the certainty of a correct prediction, forcing the model to improve on difficult-to-classify samples to lower the focal loss.
-Berman et al.
-(2018) created the Lovász loss, which is an optimizable form of the Jaccardi Index and is claimed to perform better than the CE loss due to its better categorizing of small objects and lower false negatives.
-Other losses not evaluated in this work include the Baikal loss of Gonzalez & Miikkulainen (2020), the soft cross entropy of Ilievski & Feng (2017), the sigmoid F1 of Bénédict et al.
-(2022), the work of Lee et al.
-(2021), and the MCC loss of Abhishek & Hamarneh (2021).
-In particular, the approaches of this work resemble those of Lee et al.
-(2021) and Abhishek & Hamarneh (2021), but these references did not use weighting or focal transforms, and the definitions for the differentiable F1 and MCC loss functions used in these works are slightly different from the definitions in this work.
+Ref.
+[19] created the Focal Loss, which provides a correction factor $\gamma$ that reduces the effect of increasing the certainty of a correct prediction, forcing the model to improve on difficult-to-classify samples to lower the focal loss.
+Ref.
+[20] created the Lovász loss, which is an optimizable form of the Jaccardi Index and is claimed to perform better than the CE loss due to its better categorizing of small objects and lower false negatives.
+Other losses not evaluated in this work include the Baikal loss of Ref.
+[21], the soft cross entropy of Ref.
+[22], the sigmoid F1 of Ref.
+[23], the work of Ref.
+[24], and the MCC loss of Ref.
+[25].
+In particular, the approaches of this work resemble those of Ref.
+[24] and Ref.
+[25], but these references did not use weighting or focal transforms, and the definitions for the differentiable F1 and MCC loss functions used in these works are slightly different from the definitions in this work.
 The standard formulations of evaluation metrics (such as the F1 score or MCC) are not differentiable, so they cannot be directly used as loss functions.
 However, it is possible to modify them for differentiability.
-One method is treating the true / false categories as prediction probabilities instead of binary values.
+The method chosen to add differentiability is treating the true / false categories as prediction probabilities instead of binary values.
 For example, if the model predicts $p_{t}=0.7$ for a positive (negative) sample, that sample would increase the true positives (negatives) by 0.7 and the false negatives (positives) by 0.3.
 This modification can be considered as a differentiable MCC loss function.
-The focal modification of Lin et al.
-(2020) can be added to this loss function by modifying $p_{t}$ to $(p_{t})^{\gamma}$, which leads to the focal differentiable MCC loss function.
+The focal modification of Ref.
+[19] can be added to this loss function by modifying $p_{t}$ to $(p_{t})^{\gamma}$, which leads to the focal differentiable MCC loss function.
 Finally, it is possible to add class weighting: by multiplying the true positives and false negatives by a scalar W, it is possible to prioritize predictions on the positive class (W > 1) or negative class (W < 1).
 All the above changes combined lead to the weighted focal differentiable MCC loss function (Algorithm 1).
-Algorithm 1 weighted focal differentiable MCC
-  Input: predictions $pred$ and targets $y$; hyperparameters $W$ (weight) and $\gamma$ (focal loss factor)
+We choose to optimize the MCC in particular because of its various desirable properties, especially in the context of imbalanced data [26, 27].
+Algorithm 1 (binary) weighted focal differentiable MCC
+Input: predictions predpred and targets $y$; hyperparameters $W$ (weight) and $\gamma$ (focal loss factor)
  Define the following variables: # The asterisk (*) represents multiplication.
@@
  TP = $((p_{t})^{\gamma}*y).sum()*W$
-  FN = $((1-p_{t})^{\gamma}*y).sum()*W$ $W$ is used only if y = 1, that is, if the sample is in the positive class.
+ FN = $((1-p_{t})^{\gamma}*y).sum()*W$ # $W$ is used only if y = 1, that is, if the sample is in the positive class.
  FP = $((1-p_{t})^{\gamma}*(1-y)).sum()$
@@
 return -MCC # Negative for maximization
-For the weighted CE loss, the weights for the positive class used were === {15, 20, 25, 30}.
+Declaration of competing interest
+The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
+Data availability statement
+The raw data, code used to process these data, and models underlying this article are available in a GitHub repository at https://github.com/
+PedroSeber/O-GlcNAcylation_Prediction
+S1 Reproducing the Models and Plots
+The transformer models can be recreated by downloading the datasets and running the transformer_train.py file with the appropriate flags (run python transformer_train.py --help for details).
+The RNN models can be recreated by downloading the datasets and running the ANN_train.py file (modified from the original of Ref.
+[6]) with the appropriate flags (run python ANN_train.py --help for details).
+The plots can be recreated by running the make_plot2.py file (modified from the original of Ref.
+[6]) with the appropriate flag as an input (python make_plot2.py finetune for Fig.
+1, python make_plot2.py single-layer for Fig.
+2, python make_plot2.py multilayer for Fig.
+3, python
+make_plot2.py transformer
+for Fig. S1, and python make_plot2.py
+nested
+for Fig. S2).
+S2 Using the RNN Model to Predict
+O-GlcNAcylation Sites
+Following the procedures set by Ref.
+[6], the Conda environment defining the specific packages and version numbers used in this work is available as ANN_environment.yaml on https://github.com/PedroSeber/
+O-GlcNAcylation_Prediction
+. To use our trained model, run the
+Predict.py file as python Predict.py <sequence> -t <threshold> -bs <batch_size>.
+Alternatively, create an (N+1)x1 .csv with the first row as a header (such as "Sequences") and all other N rows as the actual amino acid sequences, then run the Predict.py file as python Predict.py
+<path/to/file.csv> -t <threshold> -bs <batch_size>
+. Results are saved as a new .csv file.
+S3 List of Hyperparameters Used in This Work
+  1. 1.
+[…]
+  2. 2.
+[…]
+  3. 3.
+[…]
+  4. 4.
+    For the weighted CE loss, the positive-class weights used were {15, 20, 25, 30}.
 For the focal loss, the $\alpha$ values used were {0.9375, 0.95, 0.99, 0.999, 0.9999, 0.99999, 0.999999} and the $\gamma$ values used were {0, 1, 2, 3, 5, 10, 15, 20, 30, 45, 50}.
-For the weighted focal differentiable MCC, the weights W used were === {1, 2, 3, 4, 5, 10, 15, 20} and the $\gamma$ values used were {1, 2, 3, 3.5, 4, 4.5, 5}.
-3 Results
-3.1 Transformer Encoders Have Good Predictive Power on This Dataset, but are Inferior to RNNs
-The primary goal of this subsection is to compare transformer models with the best RNN from Seber & Braatz (2024) using the modified dataset of Wulff-Fuentes et al.
-(2021) (Section 2.1) and the weighted CE loss function.
-While our transformer models surpass models published before Seber & Braatz (2024) and achieve an F1 score equal to 24.31% and an MCC equal to 22.49% (Table 1), transformers exhibit inferior recall at the same precision level (Fig.
-1) than the RNNs from Seber & Braatz (2024), leading to inferior F1 and MCC metrics.
-It should be noted that the work of Pokharel et al.
-(2023) used an ensemble of three transformer models to predict O-GlcNAcylation and obtained considerably inferior F1 and MCC metrics than the RNN of Seber & Braatz (2024) and the transformer models in this work.
-It is possible that transformers are an ill-suited architecture for this task, potentially due to the amount of data available, which is why we train primarily RNNs in the following sections.
-Figure 1: P-R curves for O-GlcNAcylation site prediction models tested on the modified dataset of Wulff-Fuentes et al.
-(2021).
-The black dotted lines are isolines of the F1 score, as labeled on the top and right sides of the plot.
-Model “YinOYang” is from Gupta & Brunak (2002); its metrics were obtained by Seber & Braatz (2024).
-Models “RNN-#” come from Seber & Braatz (2024).
-Models “Transformer-15” come from this work.
-The number after each RNN model is the LSTM module size used; the number after each transformer model is the hidden size used; the number after the semicolon for both types of models represents the window size on each side of the central S/T.
-The numbers on the “RNN-#” and “Transformer-15” curves represent the minimum threshold for a site to be considered positive.
-“No Model” sets all sites to potentially positive – that is, sets the threshold to 0.
-Table 1: Performance metrics (in %) for the highest F1 score point of models tested on the modified dataset of Wulff-Fuentes et al.
-(2021).
-Models “Transformer-15” come from this work.
-Model “RNN-225” comes from Seber & Braatz (2024).
-The number after the RNN model is the LSTM module size used; the number after each transformer model is the hidden size used; the number after the semicolon for both types of models represents the window size on each side of the central S/T.
-Model “YinOYang” is from Gupta & Brunak (2002); its metrics were obtained by Seber & Braatz (2024).
-“No Model” sets all sites to potentially positive – that is, sets the threshold to 0.
-Metric Transformer-15; 5 win (This Work) Transformer-15; 10 win (This Work) Transformer-15; 15 win (This Work) Best Threshold 0.60 0.20 0.50 Recall (%) 20.54 26.06 26.02 Precision (%) 15.63 13.77 21.28 F1 Score (%) 17.75 18.02 23.41 MCC (%) 15.50 16.08 21.36
-Metric Transformer-15; 20 win (This Work) RNN-225; 20 win (Seber & Braatz (2024)) YinOYang (Gupta & Brunak (2002)) No Model Best Threshold 0.50 0.60 0.60 N/A Recall (%) 30.89 35.47 13.59 100 Precision (%) 20.04 36.90 8.08 2.44 F1 Score (%) 24.31 36.17 10.13 4.76 MCC (%) 22.49 34.57 7.57 0.00
-A surprising result is how the performance of the transformer models does not change significantly with the hyperparameters.
+For the weighted focal differentiable MCC, the weights W used were {1, 2, 3, 4, 5, 10, 15, 20} and the $\gamma$ values used were {1, 2, 3, 3.5, 4, 4.5, 5}.
+The transformer encoder and RNN architectures were initialized with the Kaiming uniform initialization [28] with $a=\sqrt{5}$.
+Cosine scheduling is used because it has achieved significant results with imbalanced datasets [29, 30], and it was also used by Ref.
+[6].
+S4 Supplemental Results
+S4.1 Transformer encoder models from this work surpass previous transformer models, but are inferior to RNNs
+The primary goal of this subsection is to compare transformer models with the best RNN from Ref.
+[6], labeled in this work as “RNN-225” or “RNN-225; CE”, using the modified dataset of Ref.
+[9] (Section Dataset).
+While this work’s transformer models surpass models published before Ref.
+[6] and achieve an F1 score equal to 24.31% and an MCC equal to 22.49% (Table S1), transformers exhibit inferior recall at the same precision level (Fig.
+S1) than the RNNs from Ref.
+[6], leading to inferior F1 and MCC metrics.
+It should be noted that the work of Ref.
+[11] used an ensemble of three transformer models to predict O-GlcNAcylation and obtained considerably inferior F1 and MCC metrics than the RNN of Ref.
+[6] and the transformer models in this work.
+A surprising result is how the performance of transformer encoder models does not change significantly with the hyperparameters used.
 Using large hidden sizes or adding another linear layer after the encoders slightly degrades the performance, while increasing the number of attention heads (to 4–5) and the number of stacked encoder layers (to 4) marginally improves the model.
 The most surprising result is the lack of significant improvement with increasing window size.
-As in Seber & Braatz (2024), our models perform better with larger window sizes for sizes up to 20 (Fig.
-1 and Table 1).
-The performance of a transformer encoder with a window size = 5 is only marginally inferior to the RNN from Seber & Braatz (2024) using the same window size.
+As in Ref.
+[6], our models perform better with larger window sizes for sizes up to 20 (Fig.
+S1 and Table S1).
+The performance of a transformer encoder with a window size = 5 is only marginally inferior to the RNN from Ref.
+[6] using the same window size.
 However, whereas the RNNs’ performances increase quite significantly with increasing window size, the performance of the transformers barely improves.
-The only significant performance increase comes from increasing the window size from 10 to 15, but barely any performance is gained from further increasing this window size to 20 (Fig.
-1).
-3.2 The weighted focal differentiable MCC loss function is a superior alternative to the weighted CE
+The only significant performance increase comes from increasing the window size from 10 to 15, but little performance is gained from further increasing this window size to 20 (Fig.
+S1).
 We hypothesized that changing the loss function can lead to improvements in the models’ metrics, as the weighted CE is optimizing a function that is only correlated with F1 scores and MCC.
-Four other loss functions (as described in Section 2.4) are also used to cross-validate transformer models through the same procedure used with the weighted CE function (Sections 2.2 and 3.1).
+Four other loss functions (as described in Section Loss Functions) are also used to cross-validate transformer models through the same procedure used with the weighted CE function.
 The Focal, Lovász, and differentiable F1 losses fail to achieve meaningful results.
@@
 The differentiable F1, in particular, frequently gets stuck at this local F1 maximum.
-Naturally, these models have an MCC $\approx 0$, as they are no better than random classifiers.
-Conversely, the weighted focal differentiable MCC loss is able to attain very good performance on this problem (Fig.
-2 and Table 2).
-For both architectures, its performance is superior to that of the weighted cross-entropy loss.
-As in Section 3.1, RNNs are considerably better than transformer encoders in predicting O-GlcNAcylation.
-The RNN model trained with the weighted focal differentiable MCC loss achieves an F1 score = 37.03% and an MCC = 36.58%.
-The relative and absolute increases in MCC are higher than those for the F1 score, highlighting the effectiveness of the weighted focal differentiable MCC loss, as it seeks to directly optimize the MCC.
-The weighted focal differentiable MCC loss can also be used for fine tuning; an RNN model first trained using weighted cross-entropy, then fine-tuned using the weighted focal differentiable MCC is generated (“RNN-225; CE $\rightarrow$ diff MCC”).
-This model initially followed the optimal hyperparameters (including model size) found through cross-validation using the weighted cross-entropy loss, but the learning rate and loss hyperparameters are changed to the optimal values found through cross-validation with the weighted focal differentiable MCC loss function when that loss function is used for fine-tuning.
-This fine-tuning also leads to a model that is better than the previous state-of-the-art from Seber & Braatz (2024) (Fig.
-2 and Table 2), reaching an F1 score = 36.52% and an MCC = 36.01%, values slightly lower than that of the model trained only with the weighted focal differentiable MCC loss.
-There is a significant difference between the optimal hyperparameters selected through the weighted cross-entropy and weighted focal differentiable MCC loss functions.
-Compared to the optimal values found through the weighted cross-entropy loss, the weighted focal differentiable MCC loss has an optimum at higher model hidden sizes (15 vs.
-240 for transformers; 225 vs.
-1425 for RNNs), higher batch sizes (256 vs.
-512 for transformers; 32 vs.
-128 for RNNs), lower learning rates (1×$\times 10^{-3}$ vs.
-1×$\times 10^{-4}$), and lower class weights (15 vs.
-3 for transformers; 30 vs.
-3 for RNNs).
-Figure 2: P-R curves for O-GlcNAcylation site prediction models tested on the modified dataset of Wulff-Fuentes et al.
-(2021).
-The black dotted lines are isolines of the F1 score, as labeled on the top and right sides of the plot.
-Model “RNN-225; 20 win; CE” comes from Seber & Braatz (2024).
-All others come from this work.
-The number after each RNN model is the LSTM module size used; the number after each transformer model is the hidden size used; the number after the semicolon for both types of models represents the window size on each side of the central S/T.
-The numbers on the curves represent the minimum threshold for a site to be considered positive.
-“No Model” sets all sites to potentially positive – that is, sets the threshold to 0.
-Table 2: Recall, precision, F1 score, and MCC metrics (in %) for the highest F1 score point of models tested on the modified dataset of Wulff-Fuentes et al.
-(2021).
-Model “RNN-225; CE” comes from Seber & Braatz (2024).
-All others come from this work.
-The number after each RNN model is the LSTM module size used; the number after each transformer model is the hidden size used; the number after the semicolon for both types of models represents the window size on each side of the central S/T.
-“CE” represents the weighted cross-entropy loss, and “diff MCC” represents the weighted focal differentiable MCC loss.
-“No Model” sets all sites to potentially positive – that is, sets the threshold to 0.
-Metric Transformer-15; CE (This Work) Transformer-240; diff MCC (This Work) RNN-225; CE (Seber & Braatz (2024)) Best Threshold 0.50 0.40 0.60 Recall (%) 30.89 25.70 35.47 Precision (%) 20.04 25.48 36.90 F1 Score (%) 24.31 25.59 36.17 MCC (%) 22.49 23.67 34.57
-Metric RNN-1425; diff MCC (This Work) RNN-225; CE $\rightarrow$ diff MCC (This Work) No Model Best Threshold 0.70 0.40 N/A Recall (%) 30.67 30.35 100 Precision (%) 46.70 45.84 2.44 F1 Score (%) 37.03 36.52 4.76 MCC (%) 36.58 36.01 0.00
-3.3 Stacked RNNs provide better performance, and the weighted focal differentiable MCC loss is still superior
-Seber & Braatz (2024) trained RNNs with only a single cell, a restriction that was followed by this work in Section 3.2.
-However, RNNs with multiple stacked cells may perform better than single-cell RNNs.
-To verify this hypothesis, RNN models with 2, 3, and 4 LSTM layers are tested (as per Section 2.3).
-Similarly to Section 3.2, three models are created: one trained only with the weighted CE loss, one trained only with the weighted focal differentiable MCC loss, and one first trained with the weighted CE loss, then fine-tuned with the weighted focal differentiable MCC loss.
-These models all surpass their single-RNN layer equivalents from Section 3.2.
-The model trained with the weighted CE loss (“RNN-[450,75]; CE”) displays the greatest relative improvement over its single-cell version, reaching an F1 score = 38.73 % and an MCC = 37.31%.
-Like with the single-cell RNNs, fine-tuning with the weighted focal differentiable MCC loss after training with the weighted CE leads to improvements in the model performance; this fine-tuned model (“RNN-[450,75]; CE $\rightarrow$ diff MCC”) reaches an F1 score = 38.83% and an MCC = 37.33%.
-While the performance improvement of fine-tuning is smaller in this scenario, these improved results again confirm the efficacy of fine-tuning with the weighted focal differentiable MCC loss.
-Finally, a model trained directly with the weighted focal differentiable MCC loss (“RNN-[600,75]; diff MCC”) achieves the best results, reaching an F1 score = 38.88% and an MCC = 38.20%.
-The increase in MCC is considerably higher, further corroborating the benefits of this new loss function discussed in Section 3.2.
-Overall, the “RNN-[600,75]; diff MCC” model displays a 7.5% increase in F1 score and a 10.5% increase in MCC over the “RNN-255; CE” model of Seber & Braatz (2024).
-Figure 3: P-R curves for two-layer O-GlcNAcylation site prediction models tested on the modified dataset of Wulff-Fuentes et al.
-(2021).
-The black dotted lines are isolines of the F1 score, as labeled on the top and right sides of the plot.
-The number after each RNN model is the LSTM module size used, and the number after the semicolon represents the window size on each side of the central S/T.
-The numbers on the curves represent the minimum threshold for a site to be considered positive.
-“No Model” sets all sites to potentially positive – that is, sets the threshold to 0.
-Table 3: Recall, precision, F1 score, and MCC metrics (in %) for the highest F1 score point of two-layer models tested on the modified dataset of Wulff-Fuentes et al.
-(2021).
-The number after each RNN model is the LSTM module size used, and the number after the semicolon represents the window size on each side of the central S/T.
-“CE” represents the weighted cross-entropy loss, and “diff MCC” represents the weighted focal differentiable MCC loss.
-The “No Model” (threshold = 0) baseline metrics are as in Tables 1 and 2.
-Metric RNN-[450,75]; CE (This Work) RNN-[450,75]; CE $\rightarrow$ diff MCC (This Work) RNN-[600,75]; diff MCC (This Work) Best Threshold 0.90 0.50 0.60 Recall (%) 36.69 37.62 33.11 Precision (%) 41.02 40.11 47.10 F1 Score (%) 38.73 38.83 38.88 MCC (%) 37.31 37.33 38.20
-The models with the highest average cross-validation F1 score all had only two layers and second LSTM layers with 75 neurons.
-As in Section 3.2, the optimal hyperparameters for models trained with the weighted focal differentiable MCC loss included higher model hidden sizes (600 vs.
-450 for CE), lower learning rates (5×$\times 10^{-4}$ vs.
-1×$\times 10^{-3}$), and lower class weights (2 vs.
-20).
-Surprisingly, both models had the same optimal batch size (128), and the hidden size and class weights for the model trained with the weighted CE loss moved towards those from the model trained with the weighted focal differentiable MCC loss.
-Finally, we verified that our splits for the cross-validation and test sets do not bias the models by performing five-fold nested cross-validation, following the procedure in Section 2.1, on the “RNN-[600,75]; diff MCC” model.
-The models generated by each fold had very similar performance to each other (Fig.
-A1 and Table A1), highlighting the robustness of that architecture and loss function to changes in the training/test data.
-4 Discussion
-This work generates transformer encoder (Section 3.1) and RNN (Sections 3.2 and 3.3) models to predict O-GlcNAcylation sites based on protein sequences using a modified version of an extensive public dataset Wulff-Fuentes et al.
-(2021).
-These models are also compared to the state-of-the-art RNNs trained by Seber & Braatz (2024).
-Although transformer models attain high performances on this dataset, they are inferior to the RNN models trained by Seber & Braatz (2024), reaching an F1 score of 24.31% and an MCC of 22.49% (Section 3.1, Fig.
-1, and Table 1).
-A significant difference between transformers and RNNs is that the former do not scale well with increasing window sizes, while the latter display significant improvements with larger window sizes.
-With window size = 5, both transformers and RNNs possess similar performances, but a large performance gap arises with increasing window sizes.
-Multiple reasons could be causing this inferior transformer performance, including a lack of training data.
-Although the dataset from Wulff-Fuentes et al.
-(2021) is one of the largest (if not the largest) O-GlcNAcylation datasets published, it is possible that there simply are not enough data for the transformer architecture to perform well in this task.
-Because the weighted cross-entropy loss function, which is used to train the models in Seber & Braatz (2024) and Section 3.1 of this work, only indirectly optimizes the classification performance metrics, we hypothesized that a function that could directly improve the F1 score and MCC can lead to better-performing models.
-We develop the weighted focal differentiable MCC loss to address this question (Section 2.4; Algorithm 1).
-This function allows the direct optimization of the MCC, yielding better classification models.
-We cross-validate and test transformer and RNN models using the same dataset and notice that the performance of the weighted focal differentiable MCC loss is similar to but higher than that of the weighted cross-entropy loss (Section 3.2, Fig.
-2, and Table 1).
-This loss function improves both the transformer and RNN architectures, and leads to the creation of an improved model with an F1 score = 37.03% and an MCC = 36.58%.
-Furthermore, a model first trained using the weighted cross-entropy loss, then fine-tuned using the weighted focal differentiable MCC loss achieves improved performance in this prediction task relative to a model trained using only the weighted CE loss, reaching an F1 score = 36.52% and an MCC = 36.01%.
-While these metrics are slightly smaller than the values obtained by a model directly trained with the weighted focal differentiable MCC loss, these results show how this new loss function can also be used to optimize and improve trained models at a low cost.
-The optimal hyperparameters for models trained with the weighted focal differentiable MCC loss are different from those of models trained with the weighted cross-entropy loss.
-As Seber & Braatz (2024) used only single-cell models in their work, the RNN models from Section 3.2 follow this restriction.
-Models with multiple stacked cells can have better performance, so we train RNN models with multiple LSTM cells in Section 3.3.
-These models display superior cross-validation F1 scores over models with one cell, and the best models have only 75 neurons in their 2nd cells.
-No matter the loss function used, these deeper models display better performance in the test set than the best RNN of Seber & Braatz (2024) (Fig.
-3 and Table 3).
-As in Section 3.2, fine-tuning using the weighted focal differentiable MCC loss after training with the weighted CE leads to improvements in performance over using only the weighted CE, corroborating one of the benefits of the weighted focal differentiable MCC loss.
-The model trained solely with the weighted focal differentiable MCC loss, “RNN-[600,75]; diff MCC”, achieves state-of-the-art performance in O-GlcNAcylation prediction, reaching an F1 score = 38.88% and an MCC = 38.20%.
-Overall, this is a 7.5% increase in F1 score and a 10.5% increase in MCC over the “RNN-255; CE” model of Seber & Braatz (2024).
-There are two main limitations to this work.
-On the O-GlcNAcylation prediction side, although the models developed in this work reach state-of-the-art performance in multiple independent datasets, that performance is still far away from (near-)perfect, possibly limiting the widespread usability of the models.
-It should be noted that the experimental detection of O-GlcNAcylation sites is a laborious and expensive process, so even an imperfect model may be useful.
-Having a usable model is particularly important given the relevance of O-GlcNAcylation to many diseases and the potential of O-GlcNAcylation as a target for medicines (Section 1).
-On the weighted focal differential MCC loss side, although we earnestly believe that it can improve the performance of models in almost any classification task relative to the (weighted) CE, this work tests its performance on only one task.
-Future directions primarily include testing the weighted focal differential MCC loss on multiple classification tasks to show its wide usability, including in fine-tuning models that were previously trained with the CE loss.
-Another direction is analyzing the effects of dynamically changing this loss function’s hyperparameters over the epochs to determine whether it can be further improved.
-The code used in this work is publicly available, allowing the reproduction, improvement, and reuse of this work.
-It is simple to install and run the best RNN model to predict O-GlcNAcylation sites based on the local protein sequence.
-Instructions are provided in Sections A1.1 and A1.2 of the supplemental data or the README in our GitHub <redacted for review> repository.
+These models have an MCC $\approx 0$, as they are no better than random classifiers.
+However, a transformer model trained with the weighted focal differentiable MCC was able to surpass the performance of the transformers trained with the weighted CE loss (Table S1 and Fig.
+S1), improving the F1 score to 25.59% and the MCC to 23.67%.
+Nevertheless, these metrics remain inferior to the previous state-of-the-art RNN model.
+Table S1: Performance metrics (in %) for the highest F1 score point of models tested on the modified dataset of Ref.
+[9].
+Models “Transformer-#” come from this work.
+The “RNN-225” model [6] is the previous state-of-the-art model.
+The number after the RNN model is the LSTM module size used; the number after each transformer model is the hidden size used; WS is the window size on each side of the central S/T.
+Metrics for the model “YinOYang” [7] were obtained by Ref.
+[6].
+Baseline (threshold = 0) metrics are precision = 2.44%, F1 score = 4.76%, and MCC = 0%.
+Metric Transformer-15; CE (This Work) WS = 5 WS = 10 WS = 15 WS = 20 Best Threshold 0.60 0.20 0.50 0.50 Recall (%) 20.54 26.06 26.02 30.89 Precision (%) 15.63 13.77 21.28 20.04 F1 Score (%) 17.75 18.02 23.41 24.31 MCC (%) 15.50 16.08 21.36 22.49
+Metric Transformer-240; diff MCC (This Work) RNN-225; CE (Ref.
+[6]) YinOYang (Ref.
+[7]) WS = 20 WS = 20 Best Threshold 0.40 0.60 0.60 Recall (%) 25.70 35.47 13.59 Precision (%) 25.48 36.90 8.08 F1 Score (%) 25.59 36.17 10.13 MCC (%) 23.67 34.57 7.57
+Fig S1: P-R curves for O-GlcNAcylation site prediction models tested on the modified dataset of Ref.
+[9].
+Labels, curve descriptions, and baseline metrics are as in the captions of Table S1 and Fig.
+1.
+S4.2 Nested validation of the two-layer RNN models
+Fig S2: Precision-recall curves for the “RNN-[600,75]; diff MCC” model after nested validation on the modified dataset of Ref.
+[9].
+The black dotted lines are F1 score isolines, as labeled on the top and right sides of the figure.
+Table S2: Recall, precision, F1 score, and MCC metrics (in %) for the point with highest F1 score for the “RNN-[600,75]; diff MCC” model after nested validation on the modified dataset of [9].
+Baseline metrics are as in the caption of Table S1.
+[…]
+Table S3: Recall, precision, F1 score, and MCC metrics (in %) for the point with highest F1 score for the “RNN-[450,75]; CE” model after nested validation on the modified dataset of [9].
+Baseline metrics are as in the caption of Table S1.
+[…]
+S5 Computational Resources Used
+All experiments were done in a personal computer equipped with a 13th Gen Intel® Core™ i5-13600K CPU, 64 GB of DDR4 RAM, and an NVIDIA GeForce RTX 4090 GPU.